# On the Moduli Space of Discrete Morse Functions

### v1

@@ 5.3. The space of barcodes
$Bar\coloneqq\coprod\limits_{P\text{ of barcode type}}\operatorname{Map}_{\leq}(P,\mathbb{R})$
@@
    We say that $P^{\prime}$ arises from $P^{\prime}$ by collapsing a bar if $P^{\prime}$ has one bar fewer than $P$. In that case, a collapse of a bar is an injective poset map $\eta\colon P^{\prime}\rightarrow P$. Moreover, we say that $P$ arises from $P^{\prime}$ by creation of a new bar.
@@
$\eta_{*}(\beta,\beta^{\prime})_{x}\coloneqq\begin{cases}\theta_{x}\text{ if }x\in P,\\ \theta^{\prime}_{x}\text{ if }x\notin P\end{cases}.$
@@
The euclidean edit distance on $Bar$ is defined as follows:
Let $P,P^{\prime}$ be posets of barcode type. Let $\beta_{1}\in\operatorname{Map}_{\leq}(P,\mathbb{R}),\beta_{2}\in\operatorname{Map}_{\leq}(P^{\prime},\mathbb{R})$. If $P^{\prime}$ arises from $P$ by an edit morphism $\eta\colon P\rightarrow P^{\prime}$, we define $\operatorname{d}(\beta_{1},\beta_{2})\coloneqq\operatorname{d^{euc}(\eta_{*}(\beta,\beta^{\prime}),\beta^{\prime})}$ In the general case, we define the euclidean edit distance by
@@
We topologize the space of barcodes with the euclidean edit distance.
@@
Moreover, we define the braid arrangement $\mathcal{H}(P)$ in $\operatorname{Map}_{\leq}(P,\mathbb{R})$ by $\mathcal{H}(P)\coloneqq\{H_{x}^{y}|x,y\in P\}$.
In order to define a map from the space of strict merge trees to the space of barcodes, we apply the combinatorial elder rule, see [${}^{+}$24, Definition 2.15]
@@
Let $\theta\in Mer_{<}$ be a strict merge tree with an underlying combinatorial merge tree $P$. We define a poset $\tilde{B}(\theta)$ inductively: Add for each inner node $y$ of $P$ an interval $[x,y]$ to $\tilde{B}(\theta)$, where $x$ denotes leaf of $P_{\leq y}$ with the minimal value $\theta_{x}$. For all leaf nodes $x$ that are not matched by the rule above, we add an interval $[x,y]$ to $\tilde{B}(\theta)$ where $y$ is $x$’s parent node.
@@
The map $B\colon Mer_{wb}\rightarrow Bar$ from Definition 5.28 is continuous.
@@
Let $\theta_{0}\in Mer_{wb}$ be an arbitrary well-branched merge tree and let $\epsilon>0$. We have to show that there is a $\delta>0$ such that for all $\theta\in Mer_{wb}$ with $\operatorname{d}(\theta,\theta_{0})<\delta$ we have $\operatorname{d}(B(\theta),B(\theta_{0}))<\epsilon$.
Since $\theta_{0}$ is by assumption strict, it must lie in the interior of some region of $\mathcal{O}(s(\theta))$. Thus, $\delta$ can be chosen small enough such that all $\theta$ with $\operatorname{d}(\theta,\theta_{0})<\delta$ have the same underlying combinatorial merge tree $s(\theta)$ and induce the same combinatorial barcode $\tilde{B}(\theta)$. Therefore, $B_{\lvert U_{\delta}(\theta_{0})}$ is a linear map between finite-dimensional vectorspaces and, therefore, continuous. ∎

### v2

@@ 4.4. The space of barcodes
$Bar\coloneqq\coprod\limits_{P\text{ of barcode type}}\operatorname{Map}_{\leq}(P,\mathbb{R}).$
@@
    We say that $P^{\prime}$ arises from $P$ by collapsing a bar if $P^{\prime}$ has one bar fewer than $P$. In that case, a collapse of a bar is an injective poset map $\eta\colon P^{\prime}\rightarrow P$. Moreover, we say that $P$ arises from $P^{\prime}$ by creation of a new bar.
@@
$\eta_{*}(\beta,\beta^{\prime})_{x}\coloneqq\begin{cases}\beta_{x}\text{ if }x\in P,\\ \beta^{\prime}_{x}\text{ if }x\notin P\end{cases}.$
@@
The edit path distance on $Bar$ is defined as follows:
Let $P,P^{\prime}$ be posets of barcode type. Let $\beta_{1}\in\operatorname{Map}_{\leq}(P,\mathbb{R}),\beta_{2}\in\operatorname{Map}_{\leq}(P^{\prime},\mathbb{R})$. If $P^{\prime}$ arises from $P$ by an edit morphism $\eta\colon P\rightarrow P^{\prime}$, we define $\operatorname{d}(\beta_{1},\beta_{2})\coloneqq\operatorname{d^{euc}(\eta_{*}(\beta,\beta_{2}),\beta_{2})}$ In the general case, we define the edit path distance by
@@
Proposition 4.49.
The edit path distance is a pseudo-metric on BarBar
Proof.
The proof is analogous to the one of Proposition 4.40. ∎
As for merge trees, the edit path distance on barcodes cannot be a metric due to symmetries. Again, it is possible to deal with the symmetries by using ordered barcodes to choose specific representatives.
We topologize the space of barcodes with the edit path distance.
@@
In order to define a map from the space of well-branched merge trees to the space of barcodes, we apply the combinatorial elder rule, see [CDG+24, Definition 2.15]
@@
Let $\theta\in Mer_{wb}$ be a well-branched merge tree with an underlying combinatorial merge tree $P$. We define a poset $\tilde{B}(\theta)$ inductively: Add for each inner node $y$ of $P$ an interval $[x,y]$ to $\tilde{B}(\theta)$, where $x$ denotes the leaf of $P_{\leq y}$ with the minimal value $\theta_{x}$. For all leaf nodes $x$ that are not matched by the rule above, we add an interval $[x,y]$ to $\tilde{B}(\theta)$ where $y$ is $x$’s parent node.
@@
The map $B\colon Mer_{wb}\rightarrow Bar$ from Definition 4.51 is Lipschitz continuous with constant $c=1$.
@@
Let $\theta$ and $\theta^{\prime}$ be well-branched merge trees and let $B(\theta),B(\theta^{\prime})$ be their induced barcodes. Any sequence of edit moves from $\theta$ to $\theta^{\prime}$ induces a sequence of edit moves from $B(\theta)$ to $B(\theta^{\prime})$ such that merges/splittings of inner nodes induce reorderings and additions/removals of leaves induced creations/collapses. By definition of the edit costs, additions/removals of leaves have the same cost as their induced creation/collapse of bars. Reorderings can be done simultaneously whereas splittings and merges of inner nodes need to be done consecutively, which is why the distance may decrease when proceeding to the induced barcode.
5. Future directions
The main contribution of this article is a novel framework for the space of discrete Morse functions, which turns out to naturally fit into existing related mathematical concepts. Due to this, there are several potential future directions connected to a deeper investigation of any of the mentioned connections to other mathematical concepts.
As briefly mentioned at the end of section 3, a globalization of the concept of the space of discrete Morse functions to a simple homotopy type may be useful for further studies towards the relationship to the smooth case. One could try to approach such studies in a local way, by studying paths in the space of discrete Morse functions on a CW complex $X$ and allowing them to jump to the space of discrete Morse functions on a simple homotopy equivalent CW complex $X^{\prime}$ when they reach a relevant intersection of hyperplanes, similar to the way we have treated merge trees in this work. However, it is not clear how such an approach may lead to a global understanding of the space of discrete Morse functions on a simple homotopy type. For this, it seems necessary to find a coherent way of gluing spaces of discrete Morse functions on simple homotopy equivalent complexes in a colimit-type construction. Given such a notion of space of discrete Morse functions on a simple homotopy type, one could further study the relationship between smooth and discrete Morse theory by studying an extension of Cerf’s map $\eta$ that we extended in Corollary 4.9.
In a different direction, one could further study complexes of discrete Morse matchings as sub-posets of the intersection lattice of the Morse arrangement. This may lead to a better understanding of how cancellations of critical cells affect the sublevel filtration induced by the discrete Morse function, which in turn could be used for Morse-theoretic algorithms in topological data analysis.
In a similar direction, the relationship to the spaces of merge trees and barcodes provide a novel framework for studying inverse problems in persistence.